Active-DVI (version 1.60)

The presentation tool of choice for discriminating hackers.

    See http://pauillac.inria.fr/advi/

Key bindings to write on slides (scratch writing)

Entering scratch writing mode

Press s to enter scratch writing; the cursor is modified and you must click somewhere on the page to start writing text there. Before clicking, you can

• press ? to get help,

• press ˆ G to quit scratching immediately,

• press Esc to enter the scratch writing settings mode and tune the font and font size.

Survival command kit when scratch writing

Active-DVI recognizes the following keystrokes when scratch writing on the page.

ˆ G quit  – End of scratch writing.
Esc settings – Enter the scratch writing settings mode.

In the scratch writing settings mode, the cursor is modified and you can set some charateristics of the scratch writing facility. When in doubt, press

• press ? to get help,

• press Esc to quit the setting mode.

Scratch writing settings mode keys

When in the scratch writing settings mode, the following keys have the following respective meanings:

> greater  – Increments the scratch font size.
< smaller  – Decrements the scratch font size.
b blue  – Set the color of the font to blue.
c cyan  – Set the color of the font to cyan.
g green  – Set the color of the font to green.
k black  – Set the color of the font to black.
m magenta  – Set the color of the font to magenta.
r red  – Set the color of the font to red.
w white  – Set the color of the font to white.
y yellow  – Set the color of the font to yellow.
B more blue  – Increment the blue component of the color.

G more green – Increment the green component of the cur-

rent color.

R more red – Increment the red component of the current

color.

+ positive increment – Set the color increment to positive.
− negative increment – Set the color increment to negative.
? help  – Give the list of settings available.
Esc quit  – Quit the sratch writing settings mode.

Setting the scratching font size

Just press Esc to enter the scratch writing settings mode, then > or < to increment or decrement the font size; then press Esc again, to leave the scratch writing settings mode and continue to write on the page with the new font size.

Press q to quit

Acknowledgments

Authors and contributors are : • Jun Furuse <Jun.Furuse@inria.fr>

• Pierre Weis <Pierre.Weis@inria.fr>

• Didier Rémy <Didier.Remy@inria.fr>

• Roberto Di Cosmo <Roberto.DiCosmo@inria.fr> • Xavier Leroy <Xavier.Leroy@inria.fr>

• Didier Le Botlan <Didier.Le Botlan@inria.fr> • Alan Schmitt <Alan.Schmitt@inria.fr>

• Alexandre Miquel <Alexandre.Miquel@inria.fr>

This program is distributed under the GNU LGPL. See the enclosed file COPYING.